A steel wire with mass 23.6 g and length 1.19 m is strung on a bass so that the distance from the nut to the bridge is 1.10 m. (a) Compute the linear density of the string. kg/m (b) What velocity wave on the string will produce the desired fundamental frequency of the E1 string, 41.2 Hz? in/s (c) Calculate the tension required to obtain the proper frequency. (d) Calculate the wavelength of the strings vibration (e) What is the wavelength of the sound produced in air (Assume the speed of sound in air is 343 m/s.)
Answer #1

Given that

mass m=0.0236 kg

length L=1.19 m

distance x=1.1 m

now we find the lineat density of the string

linear density \mu=[0.0236/1.19]=0.0198 kg/m

now we find the velocity

frequency =V/2x

41.2=V/2*1.1

velocity v=90.64 m/s

now we find the tension required to obtain proper frequency

tension T=90.64^2*0.0198=162.7 N

now we find the wavelength of the string's vibration

wavelength =90.64/41.2=2.2 m

now we find the wavelength of the sound produced in air

the wavelength of the sound =343/41.2=8.33 m